OpenStudy (anonymous):

The equation \[e^{-x}-x+2=0\] has one root, A. Find an integer, N, such that N

OpenStudy (anonymous):

SOLVING NUMERICALLY

OpenStudy (dumbcow):

i would use newtons method \[\large x_{n+1} =x_{n} -\frac{f(x)}{f'(x)}\] f'(x) = -e^(-x) -1 initial guess, x_0 = 1

OpenStudy (anonymous):

Wait, after the f'(x) = -e^(-x)-1, how do you work with it?

OpenStudy (dumbcow):

its a recursive process, start with x=1 then x = 1- f(1)/f'(1) then repeat with new x_value until it stops changing hence f(x)=0

OpenStudy (anonymous):

But, does e have a value, or what?

OpenStudy (dumbcow):

that is to find A. i guess to find an integer N, you could just plug in values and see when f(x) switches from pos to neg

OpenStudy (anonymous):

I'm still confused... Can you show some steps?

OpenStudy (dumbcow):

e = 2.72...

OpenStudy (dumbcow):

its just a constant

OpenStudy (anonymous):

Ok.

OpenStudy (dumbcow):

sorry i made it more complicated than it needed to be plug in 0,1,2,3 into f(x) and see what values you get, then you can guess where the zero is

OpenStudy (anonymous):

Hmm Ok. Let me try...

OpenStudy (anonymous):

I'm still stuck :(

OpenStudy (dumbcow):

f(0) = e^0 -0 +2 = 3 f(1) = e^-1 -1 +2 = 1.37 f(2) = e^-2 -2 +2 = 0.135 f(3) = e^-3 -3+2 = -0.95

OpenStudy (dumbcow):

so its decreasing and we can assume crosses x-axis somewhere between 2 and 3

OpenStudy (anonymous):

Yeah, the book says 2. Thanks :)